### v2

@@ 3.5 Analyses for SLQ with asymmetric quadrature nodes
In light of the framework of Corollary 3.3, we derive an analysis for the SLQ method in Theorem 3.16.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@
Note that the parameter bound for the number of Rademacher distributed vectors $N$ is the same in the proof [33, Section 4.3]. Thus, the focus is on the lower bound of parameters related to the error between $(m+1)$-points Gauss quadrature $I_{m}$ and Riemann Stieljes integral $I$. In [33, Section 4.3], the authors commenced with the function $\log\left(\frac{t}{\lambda_{\max}+\lambda_{\min}}\right)$ that is analytic on $[\lambda_{\min},\lambda_{\max}]$. Here, one may simply focus on the analytic function $f(t)=\log(t)$ on $[\lambda_{\min}/(\lambda_{\max}+\lambda_{\min}),\lambda_{\max}/(\lambda_{\max}+\lambda_{\min})]$. One can further find a linear transformation
@@
such that $g$ is analytic on $[-1,1]$. The semimajor axis was chosen as $\alpha_{1}=(\kappa+1)/\kappa$ so the elliptical radius is
@@
and $g(t)$ is analytically continuable on $E_{\rho_{1}}$ with foci ±$\pm 1$. Note that the maximum of $g$ on the Bernstein ellipse $E_{\rho_{1}}$ equals the one of $f$ on the scaled Bernstein ellipse $E_{\rho_{1}}^{*}$. Further, the maximum value of $f$, i.e., $M_{\rho_{1}}$, is calculated on $E_{\rho_{1}}^{*}$ with foci $1/(\kappa+1)$ and $\kappa/(\kappa+1)$
@@
with the maximum obtained at $h(-\alpha_{1})=1/2\kappa$ according to the maximum modulus theorem [4, p. 175-178]. Based on Theorem 3.12, error between $I_{m}$ and $I$ has an upper bound
@@
to guarantee the probabilistic error bound stated in Theorem 3.16. Other parameters’ bounds are kept unchanged compared with Corollary 3.3.
Next, a probabilistic relative error bound for log determinant estimation problem would be derived. We follow the framework of [33, Theorem 4.1] and propose Theorem 3.18 under the assumption $\lambda_{\max}<1$ to give straight theoretical guidance on choosing parameters for SLQ that control the relative error under certain probability. This assumption is mild since for any SPD matrix with $\lambda_{\max}>1$, it can be simply scaled to have eigenvalues $\lambda\in(0,1)$.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@
be a linear transformation and $g(t)=f(h(t))$, then $g(t)$ is analytic on $\left[-1,1\right]$. For function $f(t)=\log(t)$ that is analytic on $[\lambda_{\min},\lambda_{\max}]$, since function $g(t)$ has a singular value at $t=\mathring{t}=-(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$, let $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})<|\mathring{t}|$ as the semimajor axis length of $E_{\rho_{2}}$, with the elliptical radius
@@
then $g(t)$ is again analytically continuable in the open Bernstein ellipse $E_{\rho_{2}}$ with foci ±$\pm 1$.
In light of Theorem 3.12, one can derive
$\left|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}\right|\leq\frac{4nM_{\rho_{2}}}{(\rho_{2}^{2}-\rho_{2})\rho_{2}^{2m}},$ (17)
@@
$|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}|\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|,$ (18)
@@
With the convergence analysis for the stochastic trace estimator first proposed by [1, Section 7] and further developed in [30, Theorem 1], when $N$, the number of queries, satisfies $N\geq(24/\epsilon^{2})\cdot\log(2/\eta)$, one may obtain
@@
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon\left|\log\det A\right|\Big{\}}\geq 1-\eta.$ (19)
@@
It should be pointed out that a larger semimajor axis implies a higher convergence rate of the Lanczos quadrature [33, Section 4.1]. In Theorem 3.18, semimajor axis $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})$ is set greater than $\alpha_{1}=(\lambda_{\max}+\lambda_{\min})/\lambda_{\min}$ in Corollary 3.3 [33, Corollary 4.5]. On the other hand, $\alpha_{2}$ is smaller than $\alpha=(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$ in [33, Theorem 4.1] but the log function would reach infinity near $t=-\alpha$, so $\alpha$ is invalid in this case.
@@ 4 Further discussion
In Theorem 3.16 and Theorem 3.18, both Lanczos quadrature and Girard-Hutchinson’s trace estimator set $\epsilon/2$ as the error bound. However, one may reallocate the error tolerance so that the total number of matrix vector multiplications (MVMs) is lessened. In brief, one should try to find a pair $(\alpha,\beta)$ such that
@@
$m\geq\frac{1}{2\log(\rho_{2})}\log\left(\frac{K_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$ (20)
@@
$N\geq(6/\epsilon^{2})\cdot(\alpha/(\alpha-1))^{2}\cdot\log(2/\eta),$ (21)
where $K_{\rho_{2}}=\frac{4\alpha M_{\rho_{2}}}{\rho_{2}^{2}-\rho_{2}}$, $\rho_{2}$ and $M_{\rho_{2}}$ are defined in Theorem 3.18. Then from the perspective of minimizing MVMs with respect to $\alpha$, the problem can be viewed as
@@
$\min_{\alpha>1}\,\log\left(C\alpha\right)\left(\alpha/(\alpha-1)\right)^{2}.$ (22)
@@
$\alpha^{*}=2\log\alpha^{*}+2\log C+1,$ (23)
which is obtained through taking the derivative of the objective function and making it equal to 0. Numerical methods can be applied to solve equation (23). Theorem 4.1 summarizes the bounds of $m$ and $N$ which product is minimized through a reallocation of error.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, let $C=4M_{\rho_{2}}/(\epsilon(\rho_{2}^{2}-\rho_{2})\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right))$. If $\alpha^{*}=2\log\alpha^{*}+2\log C+1$, the Lanczos iteration step $m$ and the query number $N$ satisfy
@@ 5 Experiments
A controllable experiment is designed to compare the numbers of MVMs required to guarantee the probabilistic error bound (19) in theory based on the results of Theorem 3.16, Theorem 3.18 and Theorem 4.1. We do not compare with the theorems proposed by Cortinovis and Kressner since their analyses are based on prior knowledge about matrices (e.g., $l_{2}$ norm and stable rank of $(\log(A)-\mathrm{diag}(\log(A))$) [6]. The test is carried out on four matrices with different decaying rates, such matrices can be kernel matrices corresponding to varying smoothness [34]. We also conduct experiments on two matrices that have practical applications [7]:
@@
In the test, we set different levels of relative error $\epsilon^{*}=0.01,0.02,\ldots,0.2$, with fixed failure probability $\eta=0.1$. Note that Theorem 3.16 does not directly analyze the relative error of the log determinant approximation, but sets $\epsilon n$ as the absolute error bound. To ensure the comparability of results, $\epsilon$ in Theorem 3.16 should be replaced by
$\epsilon=\epsilon^{*}\log\det A/n.$ (24)
However, this move is somehow idealistic since the exact log determinant value in (24) is not known in advance. Hence, the numerical results of Theorem 3.16 in the comparison is biased to some extent. Despite the small defect, we may consider them as the benchmark.
For matrix $A$ with decaying rate $0.5$, Fig. 2a displays the numbers of MVMs required by different theorems to reach the varying probabilistic error bound. In terms of the other three synthetic matrices with different decaying rates, figures that depict the MVMs are much similar in shape so we would not plot them here. Fig. 2b shows that the error reallocation-based MVMs, as required by Theorem 4.1, would be smaller when eigenvalues decay slower. In Fig. 2c and Fig. 2d, one may also observe that the strategy of finding the suitable error allocation would reduce the approximations of log determinant for real matrices in the University of Florida sparse matrix collection [7].
Figure 2: Comparison of total MVMs suggested by Theorem 3.16, Theorem 3.18, Theorem 4.1 under failure probability $\eta=0.1$. (a) Results for $A\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{0.5},i=1,\ldots,5000$; (b) Results of Theorem 4.1 for $A$, $A_{1}$, $A_{2}$ and $A_{3}\in\mathbb{R}^{5000\times 5000}$ that have different decaying rates $r=0.5,1,2,3$ with eigenvalues $\lambda_{i}=0.99/i^{r},i=1,\ldots,5000$; (c) Results for matrix from nd3k in [7]; (d) Results for matrix from thermomec_TC in [7].
@@ Appendix A Jacobi Iteration Matrices have Symmetric Eigenvalue Distribution
Hence the eigenvalues of $C$ are ±$\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$, and zero repeated $(m+n-2r)$ times, where $r=\mathrm{rank}(B)$.
@@
which indicates that ±$\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$ are the eigenvalues of $C$, and zero repeated $(m+n-2r)$ times.
@@
Hence the eigenvalues of $C$ are $\gamma\pm\sigma_{1},\gamma\pm\sigma_{2},\cdots,\gamma\pm\sigma_{r}$, and $\gamma$ repeated $(m+n-2r)$ times, where $r=\mathrm{rank}(B)$.
Appendix B Tests on Theorem 3.6
@@
Details of the 4 cases can be seen in Table 1.
$A$: symmetric eigenvalue distribution? $\boldsymbol{\mu}_{1}$: symmetric absolute values?
@@
Note that the figure of $\mu(t)$ would be central symmetric about $(\bar{\lambda},\mu(\bar{\lambda}))$ if $A$ has symmetric eigenvalue distribution and $\boldsymbol{\mu}_{1}$ has symmetric absolute values. Figure 3 depicts the measure function $\mu(t)$ and the $10$ Ritz values for the three cases, which illustrates that Theorem 3.6 is valid to some extent.

### v3

@@ 3.5 Analyses for SLQ with asymmetric quadrature nodes
In light of the framework of Corollary˜3.3, we derive an analysis for the SLQ method in Theorem˜3.16.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@
Note that the parameter bound for the number of Rademacher distributed vectors $N$ is the same in the proof [33, Section 4.3]. Thus, the focus is on the lower bound of parameters related to the error between $(m+1)$-points Gauss quadrature $I_{m}$ and Riemann Stieljes integral $I$. In [33, Section 4.3], the authors commenced with the function $\log\left(\frac{t}{\lambda_{\max}+\lambda_{\min}}\right)$ that is analytic on $\lambda_{\min},\lambda_{\max}]$. Here, one may simply focus on the analytic function $f(t)=\log(t)$ on $\lambda_{\min}/(\lambda_{\max}+\lambda_{\min}),\lambda_{\max}/(\lambda_{\max}+\lambda_{\min})]$. One can further find a linear transformation
@@
such that $g$ is analytic on [1,1][-1,1]. The semimajor axis was chosen as $\alpha_{1}=(\kappa+1)/\kappa$ so the elliptical radius is
@@
and $g(t)$ is analytically continuable on $E_{\rho_{1}}$ with foci $\pm 1$. Note that the maximum of $g$ on the Bernstein ellipse $E_{\rho_{1}}$ equals the one of $f$ on the scaled Bernstein ellipse $E_{\rho_{1}}^{*}$. Further, the maximum value of $f$, i.e., $M_{\rho_{1}}$, is calculated on $E_{\rho_{1}}^{*}$ with foci $1/(\kappa+1)$ and $\kappa/(\kappa+1)$
@@
with the maximum obtained at $h(-\alpha_{1})=1/2\kappa$ according to the maximum modulus theorem [4, p. 175-178]. Based on Theorem˜3.12, error between $I_{m}$ and $I$ has an upper bound
@@
to guarantee the probabilistic error bound stated in Theorem˜3.16. Other parameters’ bounds are kept unchanged compared with Corollary˜3.3.
Next, a probabilistic relative error bound for log determinant estimation problem would be derived. We follow the framework of [33, Theorem 4.1] and propose Theorem˜3.18 under the assumption $\lambda_{\max}<1$ to give straight theoretical guidance on choosing parameters for SLQ that control the relative error under certain probability. This assumption is mild since for any SPD matrix with $\lambda_{\max}>1$, it can be simply scaled to have eigenvalues $\lambda\in(0,1)$.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@
be a linear transformation and $g(t)=f(h(t))$, then $g(t)$ is analytic on $\left[-1,1\right]$. For function $f(t)=\log(t)$ that is analytic on $\lambda_{\min},\lambda_{\max}]$, since function $g(t)$ has a singular value at $t=\mathring{t}=-(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$, let $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})<|\mathring{t}|$ as the semimajor axis length of $E_{\rho_{2}}$, with the elliptical radius
@@
then $g(t)$ is again analytically continuable in the open Bernstein ellipse $E_{\rho_{2}}$ with foci $\pm 1$.
In light of Theorem˜3.12, one can derive
(17) $\left|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}\right|\leq\frac{4nM_{\rho_{2}}}{(\rho_{2}^{2}-\rho_{2})\rho_{2}^{2m}},$
@@
(18) $|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}|\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|,$
@@
With the convergence analysis for the stochastic trace estimator first proposed by [1, Section 7] and further developed in [30, Theorem 1], when NN, the number of queries, satisfies $N\geq(24/\epsilon^{2})\cdot\log(2/\eta)$, one may obtain
@@
(19) $\mathbb{P}\Big\{\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon\left|\log\det A\right|\Big\}\geq 1-\eta.$
@@
It should be pointed out that a larger semimajor axis implies a higher convergence rate of the Lanczos quadrature [33, Section 4.1]. In Theorem˜3.18, semimajor axis $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})$ is set greater than $\alpha_{1}=(\lambda_{\max}+\lambda_{\min})/\lambda_{\min}$ in Corollary˜3.3 [33, Corollary 4.5]. On the other hand, $\alpha_{2}$ is smaller than $\alpha=(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$ in [33, Theorem 4.1] but the log function would reach infinity near $t=-\alpha$, so $\alpha$ is invalid in this case.
@@ 4 Further discussion
In Theorem˜3.16 and Theorem˜3.18, both Lanczos quadrature and Girard-Hutchinson’s trace estimator set $\epsilon/2$ as the error bound. However, one may reallocate the error tolerance so that the total number of matrix vector multiplications (MVMs) is lessened. In brief, one should try to find a pair $(\alpha,\beta)$ such that
@@
(20) $m\geq\frac{1}{2\log(\rho_{2})}\log\left(\frac{K_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$
@@
(21) $N\geq(6/\epsilon^{2})\cdot(\alpha/(\alpha-1))^{2}\cdot\log(2/\eta),$
where $K_{\rho_{2}}=\frac{4\alpha M_{\rho_{2}}}{\rho_{2}^{2}-\rho_{2}}$, $\rho_{2}$ and $M_{\rho_{2}}$ are defined in Theorem˜3.18. Then from the perspective of minimizing MVMs with respect to $\alpha$, the problem can be viewed as
@@
(22) $\min_{\alpha>1}\,\log\left(C\alpha\right)\left(\alpha/(\alpha-1)\right)^{2}.$
@@
(23) $\alpha^{*}=2\log\alpha^{*}+2\log C+1,$
which is obtained through taking the derivative of the objective function and making it equal to 0. Numerical methods can be applied to solve equation (23). Theorem˜4.1 summarizes the bounds of $m$ and $N$ which product is minimized through a reallocation of error.
@@
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, let $C=4M_{\rho_{2}}/(\epsilon(\rho_{2}^{2}-\rho_{2})\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right))$. If $\alpha^{*}=2\log\alpha^{*}+2\log C+1$, the Lanczos iteration step $m$ and the query number $N$ satisfy
@@ 5 Experiments
A controllable experiment is designed to compare the numbers of MVMs required to guarantee the probabilistic error bound (19) in theory based on the results of Theorem˜3.16, Theorem˜3.18 and Theorem˜4.1. We do not compare with the theorems proposed by Cortinovis and Kressner since their analyses are based on prior knowledge about matrices (e.g., $l_{2}$ norm and stable rank of $(\log(A)-\mathrm{diag}(\log(A))$) [6]. The test is carried out on four matrices with different decaying rates, such matrices can be kernel matrices corresponding to varying smoothness [34]. We also conduct experiments on two matrices that have practical applications [7]:
@@
In the test, we set different levels of relative error $\epsilon^{*}=0.01,0.02,\ldots,0.2$, with fixed failure probability $\eta=0.1$. Note that Theorem˜3.16 does not directly analyze the relative error of the log determinant approximation, but sets $\epsilon n$ as the absolute error bound. To ensure the comparability of results, $\epsilon$ in Theorem˜3.16 should be replaced by
(24) $\epsilon=\epsilon^{*}\log\det A/n.$
However, this move is somehow idealistic since the exact log determinant value in (24) is not known in advance. Hence, the numerical results of Theorem˜3.16 in the comparison is biased to some extent. Despite the small defect, we may consider them as the benchmark.
For matrix $A$ with decaying rate $0.5$, Fig.˜2a displays the numbers of MVMs required by different theorems to reach the varying probabilistic error bound. In terms of the other three synthetic matrices with different decaying rates, figures that depict the MVMs are much similar in shape so we would not plot them here. Fig.˜2b shows that the error reallocation-based MVMs, as required by Theorem˜4.1, would be smaller when eigenvalues decay slower. In Fig.˜2c and Fig.˜2d, one may also observe that the strategy of finding the suitable error allocation would reduce the approximations of log determinant for real matrices in the University of Florida sparse matrix collection [7].
Figure 2: Comparison of total MVMs suggested by Theorem˜3.16, Theorem˜3.18, Theorem˜4.1 under failure probability $\eta=0.1$. (a) Results for $A\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{0.5},i=1,\ldots,5000$; (b) Results of Theorem˜4.1 for AA, $A_{1}$, $A_{2}$ and $A_{3}\in\mathbb{R}^{5000\times 5000}$ that have different decaying rates $r=0.5,1,2,3$ with eigenvalues $\lambda_{i}=0.99/i^{r},i=1,\ldots,5000$; (c) Results for matrix from nd3k in [7]; (d) Results for matrix from thermomec_TC in [7].
@@ Appendix A Jacobi Iteration Matrices have Symmetric Eigenvalue Distribution
Hence the eigenvalues of $C$ are $\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$, and zero repeated (m+n2r)(m+n-2r) times, where $r=\mathrm{rank}(B)$.
@@
which indicates that $\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$ are the eigenvalues of $C$, and zero repeated (m+n2r)(m+n-2r) times.
@@
Hence the eigenvalues of $C$ are $\gamma\pm\sigma_{1},\gamma\pm\sigma_{2},\cdots,\gamma\pm\sigma_{r}$, and $\gamma$ repeated (m+n2r)(m+n-2r) times, where $r=\mathrm{rank}(B)$.
Appendix B Tests on Theorem˜3.6
@@
Details of the 4 cases can be seen in Table˜1.
AA: symmetric eigenvalue distribution? $\boldsymbol{\mu}_{1}$: symmetric absolute values?
@@
Note that the figure of $\mu(t)$ would be central symmetric about $(\bar{\lambda},\mu(\bar{\lambda}))$ if $A$ has symmetric eigenvalue distribution and $\boldsymbol{\mu}_{1}$ has symmetric absolute values. Figure 3 depicts the measure function $\mu(t)$ and the $10$ Ritz values for the three cases, which illustrates that Theorem˜3.6 is valid to some extent.